# Semantic Density: Uncertainty Quantification in Semantic Space for Large Language Models

### v2

@@ Abstract
With the widespread application of Large Language Models (LLMs) to various domains, concerns regarding the trustworthiness of LLMs in safety-critical scenarios have been raised, due to their unpredictable tendency to hallucinate and generate misinformation. Existing LLMs do not have an inherent functionality to provide the users with an uncertainty metric for each response it generates, making it difficult to evaluate trustworthiness. Although a number of works aim to develop uncertainty quantification methods for LLMs, they have fundamental limitations, such as being restricted to classification tasks, requiring additional training and data, considering only lexical instead of semantic information, and being prompt-wise but not response-wise. A new framework is proposed in this paper to address these issues. Semantic density extracts uncertainty information for each response from a probability distribution perspective in semantic space. It has no restriction on task types and is "off-the-shelf" for new models and tasks. Experiments on seven state-of-the-art LLMs, including the latest Llama 3 and Mixtral-8x22B models, on four free-form question-answering benchmarks demonstrate the superior performance and robustness of semantic density compared to prior approaches.
@@ 1 Introduction
Large language models (LLMs) have revolutionalized many domains, such as conversational agents [27], code generation [33], and mathematical discovery [31]. Given their ability for general reasoning and adaptability to new tasks, LLMs are utilized increasingly in safety-critical applications, including healthcare [34] and finance [42]. However, existing LLMs have an unpredictable tendency to hallucinate [24], leading to misleading information and risky behaviors. Responses are generated without quantitative indicators for their uncertainty, making it difficult to evaluate how trustworthy they are. As a result, concerns have been raised about their safety [46], hindering a deeper utilization of LLMs in risk-sensitive domains [4].
@@
Although significant resources have been invested in LLM development, leading to a rapid pace in new model releases, only little progress has been made in building an uncertainty quantification framework for LLMs. An ideal outcome of such a system would be a quantitative metric associated with each response that can be used as an uncertainty/confidence indicator. Users can then build on this metric to evaluate the trustworthiness of LLM responses, e.g., establish an automatic system that triggers a warning if the response uncertainty reaches a pre-defined threshold.
@@
Following this line of thought, several techniques have been proposed in the literature to extract the uncertainty/confidence score from LLMs. In addition to the baselines that directly ask the LLM itself to evaluate its own answers [36, 15], one further step was to integrate traditional uncertainty estimatition/calibration methods into LLMs [5, 43, 45]. However, due to the nature of these traditional methods, they only work on classification problems, not free-form natural language generation (NLG) tasks, which are more general and challenging. Another direction was to fine-tune the original model [19] or train an additional layer or classifier [3, 22] to output uncertainty indicators for the responses. The main drawback is that these approaches are not "off-the-shelf" for new tasks and models: additional task-specific training labels of the ground-truth confidence are needed, and the training needs to be done in a model-specific manner, limiting their applicability.
@@
Most importantly, prior work still treats LLM outputs as traditional auto-regressive predictions [23], i.e., the generated responses are simply handled as sequences of tokens/words, considering only their lexical uncertainty. However, due to the unique nature of free-form NLG, tokens that are lexically different may be semantically similar. In most LLM applications, decisions depend on the semantics of the generated responses, and the same semantics can be stated using different words or sentence structures, leading to different lexical tokens. Therefore, uncertainty in semantic space is a more essential indicator for trustworthiness of LLM responses than lexical uncertainty.
@@
Semantic entropy [16] is the state-of-the-art (SOTA) technique in semantic uncertainty [21]. However, its current design has two intrinsic limitations. First, the returned uncertainty score is prompt-wise, i.e., the semantic entropy is calculated for each prompt, instead of each response. Considering that LLMs can generate diverse responses for the same prompt, using the same uncertainty score for different responses is problematic [20]. Second, semantic entropy only evaluates semantic equivalence between the generated responses, which is a binary one-cut measurement that returns whether two responses are semantically equivalent or not. It doesn’t take into account the more fine-grained semantic differences between the responses, which encode information that can make uncertainty quantification more precise.
@@
To fill these gaps, a framework is developed in this paper for a new uncertainty metric, semantic density (SD), that can quantify the uncertainty of LLM responses in semantic space. Semantic density rebuilds the output probability distribution from a semantic perspective, and extracts an uncertainty indicator analogous to probability density. The proposed semantic density metric has the following advantages: (1) It does not need any further training or fine-tuning of the original LLM; it is an "off-the-shelf" tool that can be directly applied to any pre-trained LLMs without modifying them; (2) it does not pose any restrictions on the problem type; in particular, it works for general free-form generation tasks (3) the returned metric is response-wise, making it possible to evaluate trustworthiness of each response; and (4) it takes the fine-grained semantic differences between resposes into account, which makes uncertainty quantification more precise.
@@
The performance of the semantic density metric was compared with six existing uncertainty quantification methods designed for LLMs across four question-answering benchmark datasets. All the approaches were tested on seven SOTA LLMs, including the latest Llama 3 and Mixtral-8x22B models. Semantic density performed significantly better than the alternatives across the board, suggesting that it forms a promising foundation for evaluating trustworthiness of LLM responses.
@@ 2 Related Work
Although the main focus of LLM community is still on developing new models with better performance, a number of studies aim at measuring uncertainty in LLMs. This section summarizes their basic ideas and potential limitations, which are then targeted in the development of semantic density.
The first direction is to ask the LLM to evaluate the uncertainty of its own responses. Tian et al. [36] performed an empirical study showing that the inherent conditional probabilities are poorly calibrated in existing LLMs with RLHF (reinforcement learning from human feedback), and that verbal confidence estimates provided by the LLMs are better calibrated. Kadavath et al. [15] developed an approach where the LLM was asked to evaluate the correctness of its own answer, in the form of the probability "P(True)" that its own answer is correct. An additional “value” head was also trained to predict "P(True)", but it turned out not to generalize well to out-of-distribution datasets. In general, the performance of model self-evaluation is not as good as other more advanced uncertainty quantification methods [16].
@@
A third direction is to perform supervised learning, either by fine-tuning the original LLM or adding additional layer/classifier, to create uncertainty indicators. For example Lin et al. [19] fine-tuned the GPT-3 model to verbally express its own confidence level, and Azaria and Mitchell [3] trained an additional classifier to return the truthful probability of each generated response, based on the hidden layer activations of the LLM. Liu et al. [22] proposed the LitCab framework, in which a single linear layer is trained over the LLM’s last hidden layer representations to predict a bias term, and the model’s logits are then altered accordingly to update the response confidence. Since these methods are model-specific and need additional task-specific training labels, they cannot be readily applied to new models and tasks. In comparison, as an unsupervised method, semantic density is "off-the-shelf" for any new models and tasks, without need of additional data or modifications to the original LLMs.
@@
A common limitation of all the above methods is that they only consider the lexical information, and do not take into account the semantic relationships between responses. Yet semantics is critical in analyzing LLM outputs. As a fourth direction, semantic entropy [16] is a SOTA technique that quantifies semantic uncertainty for LLMs. It works by grouping the generated samples based on their semantic equivalence, then generating an entropy-based indicator as an uncertainty metric. Although its performances is promising compared to the other approaches, as discussed in Section 1, it has two intrinsic limitations: (1) The generated semantic entropy is prompt-wise instead of response-wise; and (2) Only a one-cut equivalence relationship is considered during semantic analysis. The proposed semantic density improves over these two aspects by providing a response-specific uncertainty indicator and analyzing the relationship among generated responses in a fine-grained semantic space.
@@
Besides the above four major directions, uncertainty quantification for LLMs has been explored from other angles as well. Lin et al. [20] tested several simple baselines, among which a straightforward measurement of semantic dispersion is robust in evaluating selective response generation. Similarly, Manakul et al. [24] proposed a framework with several variants that use sampling consistency for detecting hallucinations. These two studies assume a very restricted condition in which the original sampling likelihoods of each response are not used; thus, the uncertainty information extracted by these methods is limited. The study by Ling et al. [21] focused on a specific in-context learning setting, aiming to decompose the uncertainty of LLMs into that caused by demonstration quality and that caused by model configuration. These experiments were limited to classification problems. Similarly, Hou et al. [10] decomposed the uncertainty into data uncertainty and model uncertainty in a prompt-wise approach. Xiao and Wang [44] studied the connections between hallucination and predictive uncertainty, showing that higher uncertainty is positively correlated with higher chance to hallucinate. This study validates the importance of a reliable uncertainty measurement in detecting hallucinations of LLMs. Finally, Huang et al. [11] perform an explorative study using simple baselines on uncertainty measurement for LLMs, highlighting the need for more advanced uncertainty quantification methods developed exclusively for LLMs. This is the goal for the current paper as well.
@@ 3 Methodology
This section first defines the LLM uncertainty quantification problem, then describes the design principles and technical details of each algorithmic component, and concludes with a summary of the entire framework.
@@ 3.1 Problem Statement
Given a pre-trained LLM, an input prompt $\bm{x}$, and an output sequence $\bm{y}=[y_{1},y_{2},\cdots,y_{L}]$, where $L$ is the number of tokens in $\bm{y}$, the target is to produce an uncertainty metric that has a monotonic relationship with the probability of $\bm{y}$ to be true. Note that this metric should be response-wise, i.e., it is calculated for a specific $\bm{y}$ given $\bm{x}$. The metric can be used as a quantitative indicator for whether a specific response $\bm{y}$ can be trusted.
@@ 3.3 Semantic Density Estimator
Given the semantic space $\mathbb{S}$ defined in Section 3.2, the underlying probability distribution from which the LLM samples in $\mathbb{S}$ provides critical uncertainty information: If a response is semantically close to many highly probable samples, it should be more trustworthy compared to a response that is semantically distant from the major sampling possibilities. A classical technique for estimating probability density is kernel density estimation (KDE) [32, 28]. However, the standard KDE only works for continuous variables, whereas the LLM outputs are discrete, i.e., sequences of tokens selected from a finite vocabulary. One possible way to extend KDE to accomodate LLM outputs is to build a density estimator as
@@
where $\bm{x}$ is the input prompt, $\bm{y}_{*}$ is the target response, and $\bm{v}_{*}=E(\bm{y}_{*}|\bm{x})$. In total, $\sum_{i=1}^{M}n_{i}$ reference responses are sampled to facilitate the density estimation, where $M$ is the number of unique samples. Each $\bm{y}_{i}$ represents a unique sample; $n_{i}$ is the number of occurrences of $\bm{y}_{i}$ during sampling, $f_{i}=\frac{n_{i}}{\sum_{i=1}^{M}n_{i}}$ is the relative frequency of $\bm{y}_{i}$ during sampling, and $K(\cdot)$ is a kernel function.
@@
One advantage of the semantic density estimator of Eq. 6 is that each result $\bm{y}_{i}$ only needs to be sampled once; their relative frequency $f_{i}$ can then be estimated as $f_{i}=\frac{p(\bm{y}_{i}|\bm{x})}{\sum_{i=1}^{M}p(\bm{y}_{i}|\bm{x})}$. Given a sampling budget of $M$ reference responses, it is therefore desirable that these $M$ samples are unique (duplications will be removed before calculating Eq. 6) and have high sampling probabilities, so that they can cover more sampling regions in the semantic space. In the current implementation, diverse beam search [38], which tends to generate diverse and high probable responses, is used to sample the $M$ unique reference responses.
In practice, length-normalized probability [25, 23] is usually used to correct the length bias in sequence probability. Moreover, temperatures scaling [8, 5, 43] is a simple yet effective method for calibrating the token probabilities during sampling. Both methods can be seamlessly integrated into semantic density: The $p(\bm{y}_{i}|\bm{x})$ in Eq. 6 can be replaced with $L_{i}]{p(\bm{y}_{i}|\bm{x})}$, and the temperature changed during sampling to calibrate each $p(y_{i,j}|\cdot)$.

### v3

# Semantic Density: Uncertainty Quantification for Large Language Models through Confidence Measurement in Semantic Space

@@ Abstract
With the widespread application of Large Language Models (LLMs) to various domains, concerns regarding the trustworthiness of LLMs in safety-critical scenarios have been raised, due to their unpredictable tendency to hallucinate and generate misinformation. Existing LLMs do not have an inherent functionality to provide the users with an uncertainty/confidence metric for each response it generates, making it difficult to evaluate trustworthiness. Although several studies aim to develop uncertainty quantification methods for LLMs, they have fundamental limitations, such as being restricted to classification tasks, requiring additional training and data, considering only lexical instead of semantic information, and being prompt-wise but not response-wise. A new framework is proposed in this paper to address these issues. Semantic density extracts uncertainty/confidence information for each response from a probability distribution perspective in semantic space. It has no restriction on task types and is "off-the-shelf" for new models and tasks. Experiments on seven state-of-the-art LLMs, including the latest Llama 3 and Mixtral-8x22B models, on four free-form question-answering benchmarks demonstrate the superior performance and robustness of semantic density compared to prior approaches.
@@ 1 Introduction
Large language models (LLMs) have revolutionalized many domains, such as conversational agents [35], code generation [42], and mathematical discovery [40]. Given their ability for general reasoning and adaptability to new tasks, LLMs are utilized increasingly in safety-critical applications, including healthcare [43] and finance [51]. However, existing LLMs have an unpredictable tendency to hallucinate [31], leading to misleading information and risky behaviors. Responses are generated without quantitative indicators for their uncertainty/confidence, making it difficult to evaluate how trustworthy they are. As a result, concerns have been raised about their safety [55], hindering a deeper utilization of LLMs in risk-sensitive domains [6].
@@
Although significant resources have been invested in LLM development, leading to a rapid pace in new model releases, only little progress has been made in building an uncertainty quantification framework for LLMs. An ideal outcome of such a system would be a quantitative metric associated with each response that can be used as an uncertainty/confidence indicator. Users can then build on this metric to evaluate the trustworthiness of LLM responses, e.g., establish an automatic system that triggers a warning if the response confidence is below a pre-defined threshold.
@@
Following this line of thought, several techniques have been proposed in the literature to extract the uncertainty/confidence score from LLMs. In addition to the baselines that directly ask the LLM itself to evaluate its own answers [45, 22], one further step was to integrate traditional uncertainty estimation/calibration methods into LLMs [7, 52, 54]. However, due to the nature of these traditional methods, they only work on classification problems, not free-form natural language generation (NLG) tasks, which are more general and challenging. Another direction was to fine-tune the original model [26] or train an additional layer or classifier [4, 29] to output uncertainty/confidence indicators for the responses. The main drawback is that these approaches are not "off-the-shelf" for new tasks and models: additional task-specific training labels of the ground-truth confidence are needed, and the training needs to be done in a model-specific manner, limiting their applicability.
@@
Most importantly, prior work still treats LLM outputs as traditional auto-regressive predictions [30], i.e., the generated responses are simply handled as sequences of tokens/words, considering only their lexical uncertainty/confidence. However, due to the unique nature of free-form NLG, tokens that are lexically different may be semantically similar. In most LLM applications, decisions depend on the semantics of responses, and the same semantics can be stated using different words or sentence structures, leading to different lexical tokens. Therefore, uncertainty/confidence in semantic space is a more essential indicator for trustworthiness of LLM responses than lexical uncertainty/confidence.
@@
Semantic entropy [23] is the state-of-the-art (SOTA) technique in semantic uncertainty [28]. However, its current design has two intrinsic limitations. First, the returned uncertainty score is prompt-wise, i.e., the semantic entropy is calculated for each prompt, instead of each response. Considering that LLMs can generate diverse responses for the same prompt, using the same uncertainty score for different responses is problematic [27]. Second, semantic entropy only considers semantic equivalence, which is a binary one-cut measurement, i.e., it only returns whether two responses are considered semantically equivalent or not, without reporting how semantically different the responses are. It does not make use of the more fine-grained semantic differences among the responses, which encode information that can make uncertainty quantification more precise.
@@
To fill these gaps, a framework is developed in this paper for a new uncertainty/confidence metric, semantic density (SD), that can quantify the confidence of LLM responses in semantic space. Semantic density analyzes the output probability distribution from a semantic perspective and extracts a confidence indicator analogous to probability density. The proposed semantic density metric has the following advantages: (1) The returned confidence metric is response-wise, making it possible to evaluate the trustworthiness of each specific response; (2) it takes the fine-grained semantic differences among responses into account, which makes uncertainty quantification more precise; (3) it does not need any further training or fine-tuning of the original LLM; it is an "off-the-shelf" tool that can be directly applied to any pre-trained LLMs without modifying them; and (4) it does not pose any restrictions on the problem type; in particular, it works for general free-form generation tasks.
@@
The performance of the semantic density metric was compared with six existing uncertainty/confidence quantification methods designed for LLMs across four question-answering benchmark datasets. All the approaches were tested on seven SOTA LLMs, including the latest Llama 3 and Mixtral-8x22B models. Semantic density performed significantly better than the alternatives across the board, suggesting that it forms a promising foundation for evaluating the trustworthiness of LLM responses. The source codes for reproducing the experimental results reported in this paper are provided at: https://github.com/cognizant-ai-labs/semantic-density-paper.
@@ 2 Related Work
In the LLM literature, the terms "uncertainty" and "confidence" are used in a mixed manner. A number of studies [52, 26, 13, 5, 17] treat uncertainty and confidence as two facets of a single concept, i.e., lower confidence on one particular response corresponds to higher uncertainty (or lower certainty). Other studies try to further differentiate "uncertainty" from "confidence" [27] and only use "uncertainty" to describe the entire output distribution instead of a specific response [23, 2]. Both perspectives fall in the same research area of "uncertainty quantification/estimation" [18, 27, 10, 17], and the goals of most existing uncertainty/confidence metrics are indeed the same: to provide a quantitative indicator of the trustworthiness of LLM responses. For better coverage and clarity, this paper uses "uncertainty quantification" as a general term to describe work related to the assessment of uncertainty or confidence of LLMs, and "uncertainty/confidence" to refer to multiple metrics with mixed term definitions. The proposed semantic density is thus an indicator of response-wise "confidence".
Although the main focus of the LLM community is still on developing new models with better performance, a number of studies aim at measuring uncertainty/confidence in LLMs. This section summarizes their basic ideas and potential limitations, which are then targeted in the development of semantic density.
The first direction is to ask the LLM to evaluate the uncertainty/confidence of its own responses. Tian et al. [45] performed an empirical study showing that the inherent conditional probabilities are poorly calibrated in existing LLMs with RLHF (reinforcement learning from human feedback), and that verbal confidence estimates provided by the LLMs are better calibrated. Kadavath et al. [22] developed an approach where the LLM was asked to evaluate the correctness of its own answer, in the form of the probability "P(True)" that its own answer is correct. An additional “value” head was also trained to predict "P(True)", but it turned out not to generalize well to out-of-distribution datasets. In general, the performance of model self-evaluation is not as good as other more advanced uncertainty quantification methods [23].
@@
A third direction is to perform supervised learning, either by fine-tuning the original LLM or adding an additional layer/classifier, to create uncertainty/confidence indicators. For example Lin et al. [26] fine-tuned the GPT-3 model to verbally express its own confidence level, and Azaria and Mitchell [4] trained an additional classifier to return the truthful probability of each generated response, based on the hidden layer activations of the LLM. Liu et al. [29] proposed the LitCab framework, in which a single linear layer is trained over the LLM’s last hidden layer representations to predict a bias term, and the model’s logits are then altered accordingly to update the response confidence. Since these methods are model-specific and need additional task-specific training labels, they cannot be readily applied to new models and tasks. In comparison, as an unsupervised method, semantic density is "off-the-shelf" for any new models and tasks, without the need for additional data or modifications to the original LLMs.
@@
A common limitation of all the above methods is that they only consider the lexical information, and do not take into account the semantic relationships between responses. Yet semantics is critical in analyzing LLM outputs. As a fourth direction, semantic entropy [23] is a SOTA technique that quantifies semantic uncertainty for LLMs. It works by grouping the generated samples based on their semantic equivalence, and then generating an entropy-based indicator as an uncertainty metric. Although its performance is promising compared to the other approaches, as discussed in Section 1, it has two intrinsic limitations: (1) The generated semantic entropy is prompt-wise instead of response-wise, and (2) only a one-cut equivalence relationship is considered during semantic analysis. The second issue was considered in a recent follow-up [34] that tries to improve semantic entropy. However, it did not resolve the first issue, and the proposed uncertainty metric is still prompt-wise, limiting its utility when different responses are sampled given the same prompt. The proposed semantic density improves over these two aspects by providing a response-specific confidence indicator and analyzing the semantic relationship in a fine-grained manner.
@@
Besides the above four major directions, uncertainty quantification for LLMs has been explored from other angles as well. Lin et al. [27] tested several simple baselines, among which a straightforward measurement of semantic dispersion is robust in evaluating selective response generation. Similarly, Manakul et al. [31] proposed a framework with several variants that use sampling consistency for detecting hallucinations. These two studies assume a very restricted condition in which the original sampling likelihoods of each response are not used; thus, the uncertainty/confidence information extracted by these methods is limited. Duan et al. [8] proposes a mechanism to shift attention to more relevant components at both token and sentence levels for better uncertainty quantification. Their approach was applied to prompt-wise uncertainty metrics only. The study by Ling et al. [28] focused on a specific in-context learning setting, aiming to decompose the uncertainty of LLMs into that caused by demonstration quality and that caused by model configuration. These experiments were limited to classification problems. Similarly, Hou et al. [16] decomposed the uncertainty into data uncertainty and model uncertainty in a prompt-wise approach. Xiao and Wang [53] studied the connections between hallucination and predictive uncertainty, showing that higher uncertainty is positively correlated with a higher chance of hallucination. This study validates the importance of a reliable uncertainty measurement in detecting hallucinations of LLMs. Finally, Huang et al. [18] perform an explorative study using simple baselines on uncertainty measurement for LLMs, highlighting the need for more advanced uncertainty quantification methods developed exclusively for LLMs. This is the goal for the current paper as well.
@@ 3 Methodology
This section first defines the LLM uncertainty quantification problem, then describes the design principles and technical details of each algorithmic component, and concludes with a summary of the entire framework. The semantic space defined in Section 3.2 forms a theoretical foundation for semantic density. It can be implemented either explicitly through an embedding model or implicitly through an inference model; the latter is the approach used in this paper (details in Section 3.5).
@@ 3.1 Problem Statement
Given a pre-trained LLM, an input prompt $\bm{x}$, and an output sequence $\bm{y}=[y_{1},y_{2},\cdots,y_{L}]$, where $L$ is the number of tokens in $\bm{y}$, the target is to produce a confidence metric that is positively correlated with the probability of $\bm{y}$ to be true. Note that this metric should be response-wise, i.e., it is calculated for a specific $\bm{y}$ given $\bm{x}$. The metric can be used as a quantitative indicator for whether a specific response $\bm{y}$ can be trusted.
@@ 3.3 Semantic Density Estimator
Given the semantic space $\mathbb{S}$ defined in Section 3.2, the underlying probability distribution from which the LLM samples in $\mathbb{S}$ provides critical information: If a response is semantically close to many highly probable samples, it should be more trustworthy compared to a response that is semantically distant from the major sampling possibilities. A classical technique for estimating probability density is kernel density estimation (KDE) [41, 37]. However, the standard KDE only works for continuous variables, whereas the LLM outputs are discrete, i.e., sequences of tokens selected from a finite vocabulary. One possible way to extend KDE to accommodate LLM outputs is to build a density estimator as
@@
where $\bm{x}$ is the input prompt, $\bm{y}_{*}$ is the target response, i.e., the response that needs confidence estimation, and $\bm{v}_{*}=E(\bm{y}_{*}|\bm{x})$. In total, $\sum_{i=1}^{M}n_{i}$ reference responses are sampled to facilitate the density estimation, where $M$ is the number of unique samples. Each $\bm{y}_{i}$ represents a unique sample; $n_{i}$ is the number of occurrences of $\bm{y}_{i}$ during sampling, $f_{i}=\frac{n_{i}}{\sum_{i=1}^{M}n_{i}}$ is the relative frequency of $\bm{y}_{i}$ during sampling, and $K(\cdot)$ is a kernel function.
@@
One advantage of the semantic density estimator of Eq. 6 is that each result $\bm{y}_{i}$ only needs to be sampled once; their relative frequency $f_{i}$ can then be estimated as $f_{i}=\frac{p(\bm{y}_{i}|\bm{x})}{\sum_{i=1}^{M}p(\bm{y}_{i}|\bm{x})}$. Given a sampling budget of $M$ reference responses, it is therefore desirable that these $M$ samples are unique (duplications will be removed before calculating Eq. 6) and have high sampling probabilities, so that they can cover more sampling regions in the semantic space. In the current implementation, diverse beam search [47], which tends to generate diverse and highly probable responses, is used to sample the $M$ unique reference responses.
In practice, length-normalized probability [32, 30] is usually used to correct the length bias in sequence probability. Moreover, temperature scaling [14, 7, 52] is a simple yet effective method for calibrating the token probabilities during sampling. Both methods can be seamlessly integrated into semantic density: The $p(\bm{y}_{i}|\bm{x})$ in Eq. 6 can be replaced with $L_{i}]{p(\bm{y}_{i}|\bm{x})}$, and the temperature changed during sampling to calibrate each $p(y_{i,j}|\cdot)$.